rose12345:

c. If two points on a line are A(4, 6) and B(8, −8), the rise is __________, and the run is __________, so the slope of the line is __________.

luhivqqcherry:

rise is =-14 and run is = 4 therefore the slope of the line is -7/2.

velmalovesshaggy145:

rise=y^2-y^1= -8 - 6 =-14 Run=x^2-x^1 = 8-4 =4 divide -14 by 4

velmalovesshaggy145:

simplify

velmalovesshaggy145:

You're welcome

velmalovesshaggy145:

Slope is always Rise/Run, and if it can be simplified, simplify it

lalalalalalalalalal:

hello